OpenStudy (anonymous):

How many grams of nitrogen are in 2.00 grams of nitroglycerin C3H5N3O9? (Atomic weights N: 14.007 C: 12.0107 H: 1.00794 O: 15.9994)?

OpenStudy (anonymous):

0.18504 g

OpenStudy (anonymous):

Actually it's 0.370 g I'm just not sure how to get there

OpenStudy (anonymous):

oh....really?! ok, lemme try again

OpenStudy (anonymous):

LOL, forgot to multiply the two

OpenStudy (anonymous):

it's just simple math really, so we're looking for Nitrogen right? since we need the answer in grams, we convert everything to grams 1.01*5 = 5.05 H 16.00*9 = 144 O 14.01*3= 42.03 N 12.01*3 = 36.03 C then we just do 42.03(all the nitrogens only)/227.11(all of the atoms combined) which is equal to 0.185g but since there are TWO grams of nitroglycerin...we multiply 0.185*2= 0.307g

OpenStudy (anonymous):

@I_freak_kids_out_16

OpenStudy (anonymous):

Okay. Thank!

OpenStudy (anonymous):

No problemo